# Bipartite Fluctuations and Charge Fractionalization in Quantum Wires

### v3

@@ Abstract
We introduce a quantum information method for measuring fractional charges in ballistic quantum wires generalizing bipartite fluctuations to the chiral quasiparticles in Luttinger liquids, i.e. adding charge and current fluctuations in a region of the wire. Bipartite fluctuations at equilibrium are characterized through a logarithmic scaling with distance encoding the entangled nature of these fractional charges in one-dimensional (1D) fluids. This approach clarifies the physical meaning of the dephasing factor of electronic interferences in a ballistic ring geometry at zero temperature, as a result of charge fractionalization. We formulate an analogy towards ground-state energetics. We show how bipartite current fluctuations represent a useful tool to locate quantum phase transitions associated to Mott physics. We address a spin chain equivalence and verify the fractional charges through an algorithm such as Density Matrix Renormalization Group (DMRG). Adding a potential difference between the two sides (parties) of the wire, bipartite fluctuations can detect a bound state localized at the interface through the Jackiw-Rebbi model coexisting with fractional charges.
preprint: APS/123-QED
@@
Introduction.— In 1D quantum systems, electronic interactions give rise to unique and paradigmatic phenomena [1, 2]. The ground-state wavefunction of a Luttinger liquid has a Jastrow form similar to the Laughlin wavefunction [3, 4]. Low-dimensional ballistic quantum wires then are characterized by charge fractionalization similar to the fractional quantum Hall physics, but chiral quasiparticles can move in both directions, left or right [3, 5]. It is yet delicate to measure those fractional charges. The conductance of a quantum wire reaches $\frac{e^{2}}{h}$ as for free electrons, assuming spin-polarized electrons, which can be understood as a Fabry-Perot resonance of fractional wavepackets [5, 6, 7]. Signatures of charge fractionalization can be found through momentum-resolved tunneling when injecting an electron with fixed chirality in the middle of a wire [8, 9]; this measure reveals the difference or Asymmetry between the (fractional) charges moving in different directions which is proportional to the Luttinger parameter. Interferometry measures in a ballistic wire allows us to access informations on dephasing effects from interactions, e.g. at finite temperature the dephasing time reveals the associated electron lifetime in a Luttinger liquid [10, 11]. A dephasing time proportional to the inverse of temperature was reported in ballistic rings of GaAs [12], but both electron-electron interactions [10] and dephasing from a surrounding metallic gate [13] can explain the signal. In chiral edge states of fractional quantum Hall systems various tools are developed and proposed to measure the fractional charges with great accuracy [14] such as current noise at zero frequency [15, 16] and Ramsey interferometry in time [17].
@@
To motivate further the present work, we observe that direct measures of many-body entangled systems, e.g. through bipartite entanglement entropy [18, 19], are also showing a huge interest in the community building a bridge with quantum information theory. For Luttinger liquids, this entropy shows a logarithmic behavior with distance [20]. The 1D Fermi gas and the Luttinger liquid reveal the same entanglement entropy. In this work, our goal is to clarify how we can reveal fractional charges of 1D quantum wires from such equilibrium methods in a quantum information perspective. From this angle, it is useful to take the point of view of bipartite charge fluctuations [21] that we introduced as a possible marker of the many-body entangled nature of ground states e.g. in Luttinger liquids taking place between two macroscopic regions of the sample; this gives rise to a logarithmic behavior as a function of length or as a function of time similar to the entanglement entropy. Bipartite charge fluctuations mean that we study the variance (noise) associated to the charge or number of electrons in one of the two regions of the wire. This probe reveals the Luttinger parameter $K$ as a prefactor which has the significance of the compressibility from an analogy to statistical physics [21]. To reveal precisely the fractional charges $f_{R}=\frac{1+K}{2}$ and $f_{L}=\frac{1-K}{2}$ in units of the electron charge [3, 5, 8, 9, 10], in this work, we introduce the bipartite chiral charge fluctuations i.e. bipartite fluctuations of charge associated to the true chiral quasiparticles which move to the left and to the right. This will be obtained when adding bipartite charge and current fluctuations: Eqs. (7), (8) and (9) will reveal the fractional charges i.e. $f_{L}^{2}+f_{R}^{2}$ as a prefactor of the logarithmic scaling. From Landauer’s point of view, noise here is a quantum signal i.e. it precisely measures the entangled nature of chiral quasiparticles. We clarify the link with the dephasing of mesoscopic interferences at zero temperature in Eq. (12). Bipartite charge fluctuations of chiral quasiparticles in the fractional quantum Hall effect are also a useful tool to probe edge states [22, 23].
Luttinger liquid and fractional charges.— The Luttinger liquid takes the form of a simple quadratic string model through bosonization [1, 2]
$H=\frac{u}{2\pi}\int_{0}^{L}dx\left(\frac{1}{K}(\partial_{x}\phi)^{2}+K(\partial_{x}\theta)^{2}\right),$ (1)
where the current density takes the form $j=\frac{uK}{\pi}\partial_{x}\theta$ and the charge density reads $\rho=\rho_{0}-\frac{1}{\pi}\partial_{x}\phi$ with $\rho_{0}$ the mean density of electrons. Here, $0<K<1$ is the Luttinger parameter and $u$ the sound velocity. For free electrons $K=1$ and $u$ is equal to the Fermi velocity $v_{F}$. We can then introduce the total charge $Q$ (normalized to $e$) and total current $J$ such that $Q=\int_{0}^{L}dx\rho(x,t)=\rho_{0}L-\frac{1}{\pi}(\phi(L)-\phi(0))$ and $J=\int_{0}^{L}dxj(x,t)=\frac{uK}{\pi}(\theta(L)-\theta(0))=uK\tilde{J}$ with $uK=v_{F}$. The wire has a length $L$ and for simplicity we study spin-polarized electrons to probe the charge dynamics. On the lattice, we have the correspondences $\sum_{i=0}^{L}c^{\dagger}_{i}c_{i}=\rho_{0}L-\frac{1}{\pi}(\phi(L)-\phi(0))$ and $\sum_{i=0}^{L}\left(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i}\right)=-\frac{2i}{\pi}(\theta(L)-\theta(0))$. The Luttinger Hamiltonian corresponds to interacting fermions hopping on a lattice through adjacent sites [1, 2].
These definitions are particularly meaningful to introduce the chiral quasiparticles which propagate at the sound velocity $u$ in the fluid. In this way, we have the identifications for the charge and current densities $\rho=n_{L}+n_{R}$ and $j=u(n_{R}-n_{L})$ in terms of the chiral densities at position $x$ moving to the right $n_{R}$ and to the left $n_{L}$ respectively. We can similarly introduce $Q_{L}=\int_{0}^{L}n_{L}(x)dx$ and $Q_{R}=\int_{0}^{L}n_{R}(x)dx$. We obtain the usual identities $Q=Q_{R}+Q_{L}$ and $J=u(Q_{R}-Q_{L})$, the charge of an electron is unity. The last equality is equivalent to $\tilde{J}=\frac{J}{uK}$ within our definitions i.e. $K\tilde{J}=Q_{R}-Q_{L}$. Therefore, suppose we have a right-moving electron injected from a reservoir with the Fermi velocity; this will give rise to the identities $Q=+1$ and $J=+v_{F}=+uK$ or $\tilde{J}=+1$ and therefore to chiral fractional charges moving to the right and to the left $Q_{R}=\frac{1+K}{2}=f_{R}$ and $Q_{L}=\frac{1-K}{2}=f_{L}$. Injecting an electron with the Fermi velocity means that the reservoir can be seen as a 1D Fermi gas with a Luttinger parameter $K=1$ [5, 9]. Similarly, we can introduce a left-moving electron from a reservoir corresponding then to $J=-v_{F}=-uK$ i.e. to $\tilde{J}=-1$ corresponding to invert the formulas for $f_{L}$ and $f_{R}$ i.e. $f_{L}=\frac{1+K}{2}$ and $f_{R}=\frac{1-K}{2}$. Momentum-resolved tunneling reveals the difference of fractional charges $f_{R}-f_{L}=\pm K$ [8, 9]. Bipartite fluctuations associated to the chiral quasiparticles will reveal the prefactor $\frac{1}{K}(f_{L}^{2}+f_{R}^{2})=\frac{1}{2}(K+K^{-1})$ which is identical for electrons propagating to the left and to the right.
The Hamiltonian can be equivalently written as
$H=\frac{uK}{4\pi}\int_{0}^{L}dx\left((\partial_{x}\theta_{R}(x))^{2}+(\partial_{x}\theta_{L}(x))^{2}\right),$ (2)
where $\theta_{R,L}=\theta\mp\frac{\phi}{K}$ [3, 10]. The associate charge variables read $\phi_{R,L}=\frac{1}{2}(-\phi\pm K\theta)$ such that $\phi=(\phi_{R}+\phi_{L})$ and $\theta=(\phi_{R}-\phi_{L})/K$. We remind the commutation relations $\phi(x),\nabla\theta(y)]=i\pi\delta(y-x)$. Integrating the chiral charge densities $\partial_{x}\phi_{R,L}$, then this reveals
$\int_{0}^{L}\partial_{x}\phi_{R,L}dx=\frac{Q\pm K\tilde{J}}{2}=Q_{R,L}.$ (3)
It is also possible to introduce quasiparticle operators for the fractional charges associated to the electron operator [3] ${\cal L}_{R,L}^{f_{R},f_{L}}(x)=\exp(-if_{R,L}\theta_{R,L}(x))$ which gives a physical interpretation to the electron Green’s function and to the electron lifetime [10].
To study bipartite fluctuations we introduce the charge and current operators in region $A$, $Q_{A}=Q(x)=\int_{0}^{x}dx^{\prime}\rho(x^{\prime})=\rho_{0}x-\frac{1}{\pi}(\phi(x)-\phi(0))$ and $J_{A}=J(x)=\int_{0}^{x}dx^{\prime}j(x^{\prime})=\frac{uK}{\pi}(\theta(x)-\theta(0))=uK\tilde{J}_{A}=uK\tilde{J}(x)$.
Fractional charges and bipartite fluctuations.— The relevance of bipartite fluctuations measures can be understood from the ground-state correlation functions [2]
@@
$\langle(\theta(x)-\theta(0))^{2}\rangle$ $=$ $\frac{1}{K}{\cal F}_{1}(x)$
${\cal F}_{1}(x)$ $=$ $\frac{1}{2}\log\left(\frac{\alpha^{2}+x^{2}}{\alpha^{2}}\right),$
with $\alpha$ the lattice spacing or short-distance cutoff. These correlation functions will be precisely related to fluctuations of the charge $Q_{A}$ and current $\tilde{J}_{A}$. The function ${\cal F}_{1}$ measures the same information i.e. scaling properties as entanglement entropy at zero temperature [20]. Bipartitioning precisely means that the region $i\in[0;\frac{x}{\alpha}]$ is $A$ and the region $i\in[\frac{x}{\alpha};\frac{L}{\alpha}]$ is $B$. Bipartite charge fluctuations in a region of the wire with $x\gg\alpha$ (in the quantum limit $x\ll\beta u$ where $\beta=\frac{1}{k_{B}T}$) read [21]
@@
with $Q_{A}=Q(x)=\sum_{i=0}^{\frac{x}{\alpha}}c^{\dagger}_{i}c_{i}=\sum_{i=0}^{\frac{x}{\alpha}}n_{i}$. From the definitions above, we also have $n=n_{L}+n_{R}$ at position $i$ along the wire. In this formula, the measure is done at any time $t$ within the ground state. The symbol ${\cal F}_{Q}(x)$ can be thought of in a symmetric way i.e. ${\cal F}_{Q(x)}={\cal F}_{Q_{A}(x)}={\cal F}_{Q_{B}(L-x)}$ when the total charge $Q_{A}+Q_{B}$ is fixed. Bipartite charge fluctuations is related to the quantum Fisher information e.g. in the Kitaev p-wave superconductor [24] and Hubbard model [25]. The quantum Fisher information is measured in quantum spin chains through the dynamical spin susceptibility [26].
@@
with the corresponding current density on the lattice $j_{i}=it(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i})$ where $\tilde{j}_{i}=j_{i}/t$ for the 1D fermions, $\tilde{J}_{A}=\sum_{i\in A}\tilde{j}_{i}$ and $t$ the hopping strength (related to the hopping term); the Planck constant $\hbar=\frac{h}{2\pi}$ is set to unity. In the continuum limit, $\tilde{j}(x)=\frac{1}{K}(n_{R}(x)-n_{L}(x))$. Equivalently, ${\cal F}_{J(x)}=\langle J(x)^{2}\rangle-\langle J(x)\rangle^{2}=\frac{v_{F}^{2}}{K\pi^{2}}\ln\frac{x}{\alpha}$. The measure again is done at any time (usually, $t$ designates both the hopping amplitude and the time). Within our definitions, we can verify the identity $\frac{\partial}{\partial t}Q(x,t)=-\frac{\partial}{\partial x}J(x,t)$.
To reveal the chiral quasiparticles, we can then introduce the total charge in the region $A$ associated to left-moving and right-moving particles, $Q_{L}=\sum_{x^{\prime}\in A}n_{L}(x^{\prime})$ and $Q_{R}=\sum_{x^{\prime}\in A}n_{R}(x^{\prime})$, which are respectively related to $\phi_{L}$ and $\phi_{R}$ variables. Then, Eqs. (Bipartite Fluctuations and Charge Fractionalization in Quantum Wires) and (6) lead to
@@
This is also interpreted as $\frac{{\cal F}_{Q(x)}+{\cal F}_{K{\tilde{J}}(x)}}{4}={\cal F}_{\frac{Q(x)\pm K\tilde{J}}{2}}={\cal F}_{Q_{R}}(x)={\cal F}_{Q_{L}}(x)$ i.e. fluctuations associated to the chiral quasiparticles related to Eq. (3); this result assumes that $\langle\phi(x)\theta(0)\rangle\rightarrow cst.$ [2]. When adding Eqs. (Bipartite Fluctuations and Charge Fractionalization in Quantum Wires) and (6), then we obtain
@@
When we sum fluctuations of charge and current in the interacting electrons system this is equivalent to measure bipartite charge fluctuations associated to the chiral quasiparticles which carry a fractional charge and reveal $\frac{1}{K}(f_{L}^{2}+f_{R}^{2})=\frac{1}{2}(K+K^{-1})$ as a prefactor. The logarithmic function ${\cal F}_{1}$ can be seen as an entanglement measure of fractional charges. This can also be understood from the electron correlator in space $\langle c_{R}(x)c^{\dagger}_{R}(0)\rangle\propto\frac{1}{2\pi}\exp(-\frac{1}{K}(f_{R}^{2}+f_{L}^{2}){\cal F}_{1}(x))$, which can be verified from Eqs. (4) and standard definitions [2, 10]. It shows how an electron at $x=0$ is correlated to the position $x$ in space (at the same time) from the spread of the wavepackets at zero temperature. We also deduce
$\frac{\pi^{3}u}{4}\frac{\partial}{\partial x}({\cal F}_{Q}(x)+{\cal F}_{\tilde{J}}(x))=\frac{\pi u}{KL}(f_{R}^{2}+f_{L}^{2}).$ (10)
This is precisely the energy to add the pair of fractional charges along the quantum wire (with $x=\frac{L}{2}$). Here, we show the potential of Eq. (8) to measure bipartite fluctuations associated to mesoscopic interferences in a ring geometry. The signal detected from mesoscopic interferences can be written as the transmission probability [10] ${\cal T}=1-(T_{0}+T_{L})+\sqrt{T_{0}T_{L}}(e^{i\pi\varphi+iK_{1}-iK_{2}}+h.c.)$; $T_{0}$ and $T_{L}$ represent the transmission probabilities to enter the mesoscopic ring at $x=0$ and to escape from the ring at $x=L_{r}$, $\varphi=\Phi/\Phi_{0}$ is the magnetic flux with $\Phi_{0}$ the flux quantum. The phases $K_{1}$ and $K_{2}$ describe the dephasing from an electrical circuit with a resistance [13] and also the dephasing from interactions in the Luttinger liquid [10]. When we perform an ensemble of measures, this phase factor measures voltage fluctuations along one path $j$ of the ring through the identity $\langle e^{i{\cal K}_{j}}\rangle=e^{-\frac{\langle{\cal K}_{j}^{2}\rangle}{2}}$ [10]. The dephasing factor corresponds to voltage fluctuations integrated along a region of the ring [10]

### v4

@@ Abstract
We introduce a quantum information method for measuring fractional charges in ballistic quantum wires generalizing bipartite fluctuations to the chiral quasiparticles in Luttinger liquids, i.e. analyzing and summing charge and current fluctuations in a region of the wire. Bipartite fluctuations at equilibrium are characterized through a logarithmic scaling with distance encoding the entangled nature of these fractional charges in one-dimensional (1D) fluids. This approach clarifies the physical meaning of the dephasing factor of electronic interferences in a ballistic ring geometry at zero temperature, as a result of charge fractionalization. We formulate an analogy towards ground-state energetics. We show how bipartite current fluctuations represent a useful tool to locate quantum phase transitions associated to Mott physics. We address a spin chain equivalence and verify the fractional charges through an algorithm such as Density Matrix Renormalization Group (DMRG). Adding a potential difference between the two sides (parties) of the wire, bipartite fluctuations can detect a bound state localized at the interface through the Jackiw-Rebbi model coexisting with fractional charges.
I Introduction
@@
In 1D quantum systems, electronic interactions give rise to unique and paradigmatic phenomena [1, 2]. The ground-state wavefunction of a Luttinger liquid has a Jastrow form similar to the Laughlin wavefunction [3, 4]. Low-dimensional ballistic quantum wires then are characterized by charge fractionalization similar to the fractional quantum Hall physics, but chiral quasiparticles can move in both directions, left or right [3, 5]. It is yet delicate to measure those fractional charges. The conductance of a quantum wire reaches $\frac{e^{2}}{h}$ as for free electrons, assuming spin-polarized electrons, which can be understood as a Fabry-Perot resonance of fractional wavepackets [5, 6, 7]. Signatures of charge fractionalization can be found through momentum-resolved tunneling when injecting an electron with fixed chirality in the middle of a wire [8, 9]; this measure reveals the difference or Asymmetry between the (fractional) charges moving in different directions which is proportional to the Luttinger parameter. Interferometry measures in a ballistic wire allows us to access information on dephasing effects from interactions, e.g. at finite temperature the dephasing time reveals the associated electron lifetime in a Luttinger liquid [10, 11]. A dephasing time proportional to the inverse of temperature was reported in ballistic rings of GaAs [12], but both electron-electron interactions [10] and dephasing from a surrounding metallic gate [13] can explain the signal. In chiral edge states of fractional quantum Hall systems various tools are developed and proposed to measure the fractional charges with great accuracy [14] such as current noise at zero frequency [15, 16] and Ramsey interferometry in time [17].
@@
To motivate the present work, we emphasize that direct measures of the entangled nature of many-body systems are possible in a laboratory through quantum information theory tools such as bipartite entanglement entropy [18, 19]. For Luttinger liquids, this entropy shows a logarithmic behavior with distance [20]. The 1D Fermi gas and the Luttinger liquid reveal the same entanglement entropy. In this work, our goal is to clarify how we can reveal fractional charges of 1D quantum wires from such equilibrium methods in a quantum information perspective. It is useful to consider charge fluctuations occurring between two macroscopic subsystems. One of us introduced this tool, through the bipartite charge fluctuations [21], as a marker of the many-body entangled nature of ground states. Bipartite charge fluctuations precisely mean that we study the variance (noise) associated to the charge or number of electrons in one of the two regions of the wire. They show a logarithmic dependence on the system size or of time in analogy to the entanglement entropy. This probe reveals the Luttinger parameter $K$ as a prefactor which has the significance of the compressibility from an analogy to statistical physics [21]. To reveal the fractional charges $f_{R}=\frac{1+K}{2}$ and $f_{L}=\frac{1-K}{2}$ in units of the electron charge [3, 5, 8, 9, 10], in this work, we introduce the bipartite chiral charge fluctuations i.e. bipartite fluctuations of charge associated to the true chiral quasiparticles which move to the left and to the right. This will be obtained when adding (summing) bipartite charge and current fluctuations: Eqs. (15), (17) and (18) will reveal the fractional charges i.e. $f_{L}^{2}+f_{R}^{2}$ as a prefactor of the logarithmic scaling. We formulate a relation to ground-state energetics in Eq. (19). Also, we clarify the link with the dephasing of mesoscopic interferences at zero temperature in Eq. (21).
From Landauer’s point of view, noise here is a quantum signal i.e. it precisely measures the entangled nature of chiral quasiparticles. Bipartite charge fluctuations of chiral quasiparticles in the fractional quantum Hall effect are also a useful tool to probe edge states [22, 23]. It is also interesting to mention that noise measures are attracting attention in the community, theoretically and experimentally, related to low-dimensional quantum systems [24, 25, 26].
II Luttinger liquid and Fractional Charges
To reach these goals, in this Section, we find it useful to introduce the model and to remind a few facts associated to fractional charges in ballistic quantum wires. The Luttinger Hamiltonian corresponds to the low-energy model associated to interacting fermions hopping on a lattice through adjacent sites [1, 2].
The Luttinger liquid Hamiltonian reads
$H=\frac{u}{2\pi}\int_{0}^{L}dx\left(\frac{1}{K}(\partial_{x}\phi)^{2}+K(\partial_{x}\theta)^{2}\right).$ (1)
Here, $u$ is the sound velocity and $K$ is the Luttinger parameter such that $K=1$ for free electrons and for interacting electrons $K<1$. The charge density reads $\rho=\rho_{0}-\frac{1}{\pi}\partial_{x}\phi$, with $\rho_{0}$ the mean density of electrons, and the current density takes the form $j=\frac{uK}{\pi}\partial_{x}\theta$. We remind the correlation relations $\phi(x),\nabla\theta(y)]=i\pi\delta(x-y)$. The correspondence with the free electron gas can be reached when modifying $\phi\rightarrow\sqrt{K}\tilde{\phi}$ and $\theta\rightarrow\tilde{\theta}/\sqrt{K}$. The ground state correlation functions of the modified fields take the symmetric forms
$\langle(\tilde{\phi}(x)-\tilde{\phi}(0))^{2}\rangle={\cal F}_{1}(x)$ (2)
$\langle(\tilde{\theta}(x)-\tilde{\theta}(0))^{2}\rangle={\cal F}_{1}(x)$
with the logarithmic function
${\cal F}_{1}(x)=\frac{1}{2}\log\left(\frac{\alpha^{2}+x^{2}}{\alpha^{2}}\right),$ (3)
and the ground state correlation functions of the original fields then take the form [2]
@@ II Luttinger liquid and Fractional Charges
$\langle(\theta(x)-\theta(0))^{2}\rangle=\frac{1}{K}{\cal F}_{1}(x).$
Here, $\alpha$ represents the lattice spacing or short-distance cutoff. The electron operator for a right or left mover reads [2]
$c_{R/L}(x)=\frac{1}{\sqrt{2\pi\alpha}}e^{i(\mp\phi+\theta)}.$ (5)
Right and Left movers refer to the direction of propagation along the wire. From results above, then we reveal the general form for the electron Green’s function
$\langle c_{R}(x)c^{\dagger}_{R}(0)\rangle=\frac{1}{2\pi\alpha}e^{-\frac{1}{K}\left(f_{L}^{2}+f_{R}^{2}\right){\cal F}_{1}(x)}.$ (6)
To understand the physical meaning of the fractional charges $f_{L}$ and $f_{R}$ in this equation, we remind here that the Luttinger liquid Hamiltonian can be equivalently written as [3, 10]
$H=\frac{uK}{4\pi}\int_{0}^{L}dx\left((\partial_{x}\theta_{R})^{2}+(\partial_{x}\theta_{L})^{2}\right),$ (7)
where $\theta_{R,L}=\theta\mp\frac{\phi}{K}$. The fractionalization mechanism can then be understood simply saying that these fields associated to left and right modes in the wire are different than the fields $\mp\phi+\theta$ defining the electron operators in Eq. (5). Therefore, a right or left-moving electron will decompose itself into fractional quasiparticles. To reveal these fractional quasiparticles in Eq. (10) we proceed as follows. We introduce the associated charge variables $\phi_{R,L}=\frac{1}{2}(-\phi\pm K\theta)$ such that $\phi=-(\phi_{R}+\phi_{L})$ and $\theta=(\phi_{R}-\phi_{L})/K$. They satisfy the commutation relations $\theta_{R/L}(x),\partial_{y}\phi_{R/L}(y)]=i\pi\delta(x-y)$. Integrating the chiral charge densities $\frac{1}{\pi}\partial_{x}\phi_{R,L}$ then we precisely reveal the selection rules for the fractional charges:
$\frac{1}{\pi}\int_{0}^{L}\partial_{x}\phi_{R,L}dx=\frac{Q\pm K\tilde{J}}{2}=Q_{R,L}.$ (8)
We have introduced the (integrated) charge $Q=-\frac{1}{\pi}\int_{0}^{L}\partial_{x}\phi(x)dx$ and the current $\tilde{J}=\frac{1}{\pi}\int_{0}^{L}\partial_{x}\theta dx$. It is then useful to introduce the integrated chiral charges such that $Q_{L}+Q_{R}=Q$ and $J=u(Q_{R}-Q_{L})=uK\tilde{J}$. We have the identifications for the charge $Q$ and current $\tilde{J}$:
$Q(L)$ $=$ $-\frac{1}{\pi}(\phi(L)-\phi(0))$
$\tilde{J}(L)$ $=$ $\frac{1}{\pi}(\theta(L)-\theta(0)).$ (9)
If we add an electron e.g. from a reservoir such that $Q=+1$ and $J=uK$ with the identity $uK=v_{F}$ for one wire [2] i.e. $\tilde{J}=+1$ then we verify the occurrence of fractional charges $Q_{R}=\frac{1+K}{2}=f_{R}$ and $Q_{L}=\frac{1-K}{2}=f_{L}$ [5, 3, 9]. In this way, the electron operator can be equivalently introduced as [10]
$\psi_{R}(x)=\frac{1}{\sqrt{2\pi\alpha}}{\cal L}_{R}^{f_{R}}\cdot{\cal L}_{L}^{f_{L}}$ (10)
where
${\cal L}_{R,L}^{f_{R},f_{L}}=e^{if_{R,L}\theta_{R,L}}.$ (11)
We emphasize that from momentum-resolved tunneling, it is possible to measure the difference between the fractional charges, $f_{R}-f_{L}=K$, from DC current or transport measures [8, 9]. The electron Green’s function in real space is not accessible in an experiment, therefore in Sec. III we provide several complementary views of Eq. (6) associated to charge and current fluctuations which may lead to further practical applications, e.g. as an interpretation of dephasing at zero temperature in a quantum wire through the physical meaning of entangled fractional charges. Related to this aim, in Sec. III we propose below to measure the factor $\frac{1}{K}\left(f_{L}^{2}+f_{R}^{2}\right){\cal F}_{1}(x)$ through bipartite charge and current fluctuations within the ground state at equilibrium. For free electrons, the function ${\cal F}_{1}$ precisely enters in the formula for bipartite charge fluctuations which acquires a nice interpretation in terms of entangled measure between two regions of the quantum fluid, a region $A$ of length $x$ and the region $B$ of length LxL-x [21]. Therefore, for interacting electrons, as we develop below, the factor $\frac{1}{K}\left(f_{L}^{2}+f_{R}^{2}\right){\cal F}_{1}(x)$ will acquire a clear understanding in terms of entangled fractional charges measurable through the dephasing profile of mesoscopic interferences in 1D ballistic rings at zero temperature i.e. in the quantum limit.
In Sec. IV we show through DMRG that bipartite current fluctuations are precisely measurable numerically e.g. in quantum spin chain analogues. In this way, it allows us to reveal the presence of fractional charges in 1D quantum fluids. It is also a very precise tool to locate quantum phase transitions e.g. associated to Mott physics. In Sec. V, we elaborate on the idea that bipartite current fluctuations can probe the presence of a localized bound state at a (topological) interface through the Jackiw-Rebbi model coexisting with fractional charges.
III Fractional Charges, Bipartite Fluctuations and Dephasing in a Ring
From the identifications between integrated charge and current in a domain of size $x$, which can be understood as $Q(x)=-\frac{1}{\pi}(\phi(x)-\phi(0))$ and $\tilde{J}(x)=\frac{1}{\pi}(\theta(x)-\theta(0))$, we can evaluate bipartite charge and current fluctuations associated to this region. We emphasize here that this region is referred to as region $A$ and the complementary region of length LxL-x will be e.g. the region $B$.
Bipartite charge fluctuations in the region of the wire with $x\gg\alpha$ (in the quantum limit $x\ll\beta u$ where $\beta=\frac{1}{k_{B}T}$) read [21]
@@ III Fractional Charges, Bipartite Fluctuations and Dephasing in a Ring
with
$Q_{A}=Q(x)=\sum_{i=0}^{\frac{x}{\alpha}}c^{\dagger}_{i}c_{i}=\sum_{i=0}^{\frac{x}{\alpha}}n_{i}.$ (13)
We introduce the electron creation (annihilation) operator on the lattice $c^{\dagger}_{i}$ $(c_{i})$ and the corresponding density $n_{i}=c^{\dagger}_{i}c_{i}$. From the definitions above, we also have $n=n_{L}+n_{R}$ at position $i$ along the wire with $n_{L}$ and $n_{R}$ referring to the density of electrons moving to the left and to the right. In this formula, the measure is done at any time $t$ within the ground state. In the continuum limit, we equivalently have the relations for the charge and current densities $\rho(x)=n_{L}(x)+n_{R}(x)$ and $j(x)=u(n_{R}(x)-n_{L}(x))$. The symbol ${\cal F}_{Q}(x)$ can be thought of in a symmetric way i.e. ${\cal F}_{Q(x)}={\cal F}_{Q_{A}(x)}={\cal F}_{Q_{B}(L-x)}$ when the total charge $Q_{A}+Q_{B}$ is fixed. Bipartite charge fluctuations is related to the quantum Fisher information e.g. in the Kitaev p-wave superconductor [27] and Hubbard model [28]. The quantum Fisher information is measured in quantum spin chains through the dynamical spin susceptibility [29].
@@
with the corresponding current density on the lattice $j_{i}=it(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i})$ where $\tilde{j}_{i}=j_{i}/t$ for the 1D fermions, $\tilde{J}(x)=\tilde{J}_{A}=\sum_{i\in A}\tilde{j}_{i}$ and $t$ the hopping strength (related to the hopping term); the Planck constant $\hbar=\frac{h}{2\pi}$ is set to unity. Hopping of electrons on the lattice results in an Hamiltonian $H=-t\sum_{i}(c^{\dagger}_{i}c_{i+1}+h.c.)$ and the current density is obtained from the Heisenberg equation of motion $\frac{d}{dt}n_{i}=i[H,n_{i}]$. The interaction term commutes with the density operator. In the continuum limit, $\tilde{j}(x)=\frac{1}{K}(n_{R}(x)-n_{L}(x))$. Equivalently, ${\cal F}_{J(x)}=\langle J(x)^{2}\rangle-\langle J(x)\rangle^{2}=\frac{v_{F}^{2}}{K\pi^{2}}\ln\frac{x}{\alpha}$. The measure again is done at any time (usually, $t$ designates both the hopping amplitude and the time). Within our definitions, we can verify the identity $\frac{\partial}{\partial t}Q(x,t)=-\frac{\partial}{\partial x}J(x,t)$.
To reveal the chiral quasiparticles, we can then introduce the total charge in the region $A$ associated to left-moving and right-moving particles, $Q_{L}=\sum_{x^{\prime}\in A}n_{L}(x^{\prime})$ and $Q_{R}=\sum_{x^{\prime}\in A}n_{R}(x^{\prime})$, which are respectively related to $\phi_{L}$ and $\phi_{R}$ variables. Then, Eqs. (III) and (14) lead to
@@
This is also interpreted as
$\frac{{\cal F}_{Q(x)}+{\cal F}_{K{\tilde{J}}(x)}}{4}={\cal F}_{\frac{Q(x)\pm K\tilde{J}}{2}}={\cal F}_{Q_{R}}(x)={\cal F}_{Q_{L}}(x).$ (16)
From Eq. (8), the chiral charge quasiparticles operators are $Q_{R/L}=\frac{(Q\pm K\tilde{J})}{2}$ such that the equation above refers to the bipartite fluctuations associated to those fractional charges. Fluctuations associated to the chiral quasiparticles in Eq. (8) then reveals the Luttinger parameter $K$. The identity $\frac{{\cal F}_{Q(x)}+{\cal F}_{K{\tilde{J}}(x)}}{4}={\cal F}_{\frac{Q(x)\pm K\tilde{J}}{2}}$ assumes that $\langle\phi(x)\theta(0)\rangle\rightarrow constant$ when $x\gg\alpha$ [2]. To precisely reveal the presence of fractional charges through $f_{L}$ and $f_{R}$, we propose to add charge and current bipartite fluctuations associated to the system of interacting electrons. From Eqs. (III) and (14), then we obtain
@@
The existence of fractional charges in Sec. II was derived from equivalences through charge and current conservations in a wire [5, 3, 9]. Therefore, to reveal the presence of those charges in the fluctuations, we show here that this also requires to sum charge and current bipartite fluctuations in a domain of the wire. When we sum these fluctuations of charge and current for the interacting electrons system this is also equivalent to measure bipartite charge fluctuations associated to the two species (colors, L or R) of chiral quasiparticles which carry a fractional charge and reveal $\frac{1}{K}(f_{L}^{2}+f_{R}^{2})=\frac{1}{2}(K+K^{-1})$ as a prefactor. The logarithmic function ${\cal F}_{1}$ then can be seen as an entanglement measure of fractional charges. This is then another way to interpret the profile of the Green’s function of the electron operator in Eq. (6) in space, that will be important to understand the presence of dephasing at zero temperature in a ballistic ring hereafter. We also deduce
$\frac{\pi^{3}u}{4}\frac{\partial}{\partial x}({\cal F}_{Q}(x)+{\cal F}_{\tilde{J}}(x))$ $=$ $\frac{\pi u}{2K}\frac{1}{x}(f_{L}^{2}+f_{R}^{2})\mid_{x=\frac{L}{2}}$ (19)
$=$ $\frac{\pi u}{KL}(f_{R}^{2}+f_{L}^{2}).$
This is precisely the energy to add the pair of fractional charges in the middle of the quantum wire (with $x=\frac{L}{2}$).
Here, we show the potential of Eq. (17) to measure bipartite fluctuations associated to mesoscopic interferences in a ring geometry. The signal detected from mesoscopic interferences can be written as the transmission probability [10] ${\cal T}=1-(T_{0}+T_{L})+\sqrt{T_{0}T_{L}}(e^{i\pi\varphi+iK_{1}-iK_{2}}+h.c.)$; $T_{0}$ and $T_{L}$ represent the transmission probabilities to enter the mesoscopic ring at $x=0$ and to escape from the ring at $x=L_{r}$, $\varphi=\Phi/\Phi_{0}$ is the magnetic flux with $\Phi_{0}$ the flux quantum. The phases $K_{1}$ and $K_{2}$ describe the dephasing from an electrical circuit with a resistance [13] and also the dephasing from interactions in the Luttinger liquid [10]. When we perform an ensemble of measures, this phase factor measures voltage fluctuations along one path $j$ of the ring through the identity $\langle e^{i{\cal K}_{j}}\rangle=e^{-\frac{\langle{\cal K}_{j}^{2}\rangle}{2}}$ [10]. The dephasing factor corresponds to voltage fluctuations integrated along a region of the ring [10]